### v2

@@ 5.1 AdS/dS
Figure 13: The Penrose diagram of dS/dS correspondence. The black and blue dashed lines represent the codimension-one hypersurfaces at $t=0$ and $w=\frac{\pi}{2}$, respectively. The black curve represents the spacelike hypersurface at $t=t_{0}$, and the orange and blue curves represent the dS slices at $w=w_{c}$, where the dual field theory is located. The intersection of these two hypersurfaces corresponds to the critical case that saturates the strong subadditivity inequality, where $\sin w_{c}\cosh t_{0}=1$.
By substituting eq. (114) into the left-hand side of the boosted strong subadditivity inequality (116), we obtain
$\text{BSSA}(\Delta\phi)=\frac{c_{\rm dS}}{48}\frac{\sin^{2}\Delta\phi}{\sin^{6}\frac{\Delta\phi}{2}\sin^{2}w_{c}\left(\frac{1}{\sin^{2}\frac{\Delta\phi}{2}\sin^{2}w_{c}}-1\right)^{\frac{3}{2}}}>0,$ (117)
which indicates that, in this case, the dual field theory remains nonlocal. In particular, at the ultraviolet boundary where $w_{c}=\frac{\pi}{2}$, there is a significant violation of the boosted strong subadditivity inequality, as indicated by the purple line in figure 14.
Figure 14: The boosted strong subadditivity inequality violate in the dS/dS holography.
@@ 5.3 The half-dS holography
By introducing a timelike boundary $\Theta=\Theta_{0}$ in dS global spacetime, the half-dS holographic model was obtained in ref. Kawamoto:2023nki . The dual field theory resides on the timelike boundary, and the bulk spacetime is confined to the region where $0<\Theta<\Theta_{0}$. The induced metric on the boundary is given by
@@
where the coordinate transformation between $\tilde{t}$ and $t$ is
$\tilde{t}=\frac{1}{\sin\Theta_{0}}\ln\frac{\cot\frac{t}{2}+1}{\cot\frac{t}{2}-1}.$ (120)
Mapping from the cylinder to the plane as in eq. (101), the metric transforms into a conformally flat form
@@
Therefore, the entanglement entropy is
$S=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{1}{(\cosh t_{0}\sin\Theta_{0}\sin\Phi_{a})^{2}}-1}}\right).$ (127)
@@
$S=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{1}{\sin^{2}\frac{\Delta\Phi}{2}}-1}}\right)=\frac{c_{\rm dS}}{6}\Delta\Phi.$ (128)
By replacing the field theory parameter $c_{\rm dS}$ with the gravitational parameter $\frac{3\ell_{\rm dS}}{2G}$, we find that the above result matches the holographic calculation presented in ref. Kawamoto:2023nki , confirming the validity of the half-dS holography. Additionally, the requirement that the effective radius of the sphere be smaller than the cosmological horizon aligns precisely with the condition given in eq. (3.6) of ref. Kawamoto:2023nki to ensure the existence of spacelike geodesics:
$r_{\text{eff}}=\ell_{\rm dS}\cosh t_{0}\sin\Theta_{0}\sin\frac{\Delta\Phi}{2}\leq\ell_{\rm dS}.$ (129)
In ref. Kawamoto:2023nki , the dual field theory in the half-dS holography was proposed to exhibit non-locality. It was concluded that when $\cosh t_{0}\leq\frac{1}{\sin\Theta_{0}}$, the entanglement entropy satisfies the strong subadditivity inequality, whereas when $\cosh t_{0}>\frac{1}{\sin\Theta_{0}}$, the strong subadditivity inequality is violated. As shown in figure 15, the orange line segment represents the region where the strong subadditivity inequality is violated, while the blue line segment represents the region where the strong subadditivity inequality is satisfied. Next, we demonstrate that even in the blue region where the strong subadditivity inequality is satisfied, the boosted strong subadditivity inequality is still violated.
Figure 15: The half-dS holography. The dual field theory resides on a timelike boundary at $\Theta=\Theta_{0}$, while the bulk spacetime covers the region $0<\Theta<\Theta_{0}$. Here we set $\cos T=\frac{1}{\cosh t}$. The blue intervals indicate $\cosh t\leq\frac{1}{\sin\Theta_{0}}$, i.e., $\Theta_{0}-\frac{\pi}{2}\leq T\leq\frac{\pi}{2}-\Theta_{0}$.
By observing eqs. (112) and (127), we find that the results from calculating the spatial interval entanglement entropy in the dS/dS holography and half-dS holography are similar. However, due to the difference in boundary-induced metrics, the boosted strong subadditivity inequality receives corresponding corrections. In the half-dS holography, the infinitesimal form of the boosted strong subadditivity inequality is
@@
where $\Delta\tilde{\Phi}=\sin\Theta_{0}\Delta\Phi$. Substituting the result for the entanglement entropy into the above expression, we can get
$\text{BSSA}(\Delta\phi)=\frac{c_{\rm dS}}{48}\frac{\sin^{2}(\Delta\phi\sin\Theta_{0})}{\sin^{6}(\frac{1}{2}\Delta\phi\sin\Theta_{0})(\frac{1}{\sin^{2}\Theta_{0}\sin^{2}(\frac{1}{2}\Delta\phi\sin\Theta_{0})}-1)^{\frac{3}{2}}}>0.$ (131)
The result is illustrated in figure 16, which indicates that, in the half-dS holography, the boundary field theory still does not satisfy the boosted strong subadditivity inequality.
Figure 16: The boosted strong subadditivity inequality violates in the half-dS holography.
We have calculated the entanglement entropy of subsystem $A$ using the generalized spherical partition function method within the dS/dS correspondence and half-dS holography frameworks. Furthermore, we have reaffirmed the conclusions from Section 3.2 by applying the boosted strong subadditivity inequality and the strong subadditivity inequality: in dS3 spacetime, when the effective radius of the sphere is smaller than the cosmological horizon, the dual field theory exhibits non-locality.
@@ 6 Conclusion and Discussion
This paper investigated the effects of the $T\bar{T}$ deformation on various holographic dualities, including AdS/CFT, dS/CFT, and dS/dS, by analyzing the entanglement entropy of the boundary field theory. When the boundary is spherical, and the subsystem is a semicircular arc, the entanglement entropy can be derived from the classical partition function on the $n$-replica manifold, with the partition function governed by the trace flow equation.
To study the entanglement entropy in the dS case, we first derived the trace flow equation for the dS/CFT by applying a double Wick rotation to the AdS case. Then, we considered a simple scenario where the subsystem is a semicircular arc on the sphere, which corresponds to global coordinates in the dS/CFT, as well as the $t=0$ moment in both AdS/dS and dS/dS. Our findings are as follows:
  • In dS/CFT, the entanglement entropy is complex, indicating that the dual field theory is non-unitary.
  • In dS/dS, the entanglement entropy is real but violates the strong subadditivity inequality, suggesting that the dual field theory is unitary but non-local.
We then examined a more complicated scenario where the boundary-induced metric is time-dependent, covering AdS/dS, dS/dS, and half-dS holography. In these cases, since the expressions for the strong subadditivity inequality and the boosted strong subadditivity inequality depend on the specific form of the metric, and the boundary-induced metric is time-dependent, the locality of the boundary field theory may vary across different time periods, as discussed in ref. Kawamoto:2023nki . To address this, we further generalized the replica method on the sphere to accommodate more general manifolds and subsystems. We first verified this method in the global coordinates of AdS/CFT, dS/CFT, and the planar BTZ case, where the results are consistent with existing perturbative calculations Chen:2018eqk ; Chen:2023eic . We then applies it to the cases where the boundary-induced metrics are time-dependent. Our results show that while the subsystem’s entanglement entropy may satisfy the strong subadditivity inequality at certain time periods in dS/dS and half-dS holographies, the boosted strong subadditivity inequality is consistently violated, indicating that the dual field theories remain non-local throughout.
@@ Acknowledgements.
We would like to thank Shan-Ming Ruan for reading our draft and providing comments. We also thank Yang He, and Shan-Ping Wu for their helpful discussion. This work was supported by the National Natural Science Foundation of China (Grants No. 12475056, No. 12235016, No. 12075101, No. 12475053, and No. 12247101), the 111 Project under (Grant No. B20063), the Major Science and Technology Projects of Gansu Province, and Lanzhou City’s scientific research funding subsidy to Lanzhou University. S.H. would appreciate the financial support from the Fundamental Research Funds for the Central Universities and Max Planck Partner Group and the Natural Science Foundation of China (NSFC) Grants No. 12075101 and No. 12235016. L.Z. is supported by the Science and Technology Development Plan Project of Jilin Province, China (No. 20240101326JC).
Appendix A The gravity calculation
A.1 The dS/Poincaré case
@@
Finally we can get the geodesic distance is
$D_{12}=\ell_{\rm dS}\arccos\left(\ell_{\rm dS}^{2}\left(1-\frac{2X_{a}^{2}}{\eta_{c}^{2}}\right)\right)\simeq-2i\ell_{\rm dS}\log\frac{2X_{a}}{\eta_{c}}+i\ell_{\rm dS}\frac{\eta_{c}^{2}}{2X_{a}^{2}}+\pi\ell_{\rm dS}.$ (137)

### v3

@@ 5.1.2 The case of $t_{0}\neq 0$
Figure 16: The Penrose diagram of dS/dS holography. The black and blue dashed lines represent the codimension-one hypersurfaces at $t=0$ and $w=\frac{\pi}{2}$, respectively. The black curve represents the spacelike hypersurface at $t=t_{0}$, and the orange and blue curves represent the dS slices at $w=w_{c}$, where the dual field theory is located. The intersection of these two hypersurfaces corresponds to the critical case that saturates the SSA, where $\sin w_{c}\cosh t_{0}=1$.
For the two cases described above, the critical positions are illustrated in figure 16. The orange region corresponds to $\sin w_{c}\cosh t_{0}>1$, where the dual field theory is non-local. The blue region corresponds to $\sin w_{c}\cosh t_{0}<1$, where the dual field theory satisfies the SSA, but it still violates the BSSA. This demonstrates that the dual field theory still exhibits non-locality in the blue region in figure 16.
@@ 5.2 The half-dS holography
For the half-dS holography, the induced metric on the boundary is given by
@@
where the coordinate transformation between $\tilde{t}$ and $t$ is $\tilde{t}=\frac{1}{\sin\theta_{0}}\ln\frac{\cot\frac{t}{2}+1}{\cot\frac{t}{2}-1}$. Mapping from the cylinder to the plane as in eq. (89), the metric transforms into a conformally flat form
@@
In order to ensure the existence of spacelike geodesics, the ref. Kawamoto:2023nki requires  [^7] Equation (3.6) of ref. Kawamoto:2023nki with the substitution $\phi_{1}-\phi_{2}=2\phi_{a}=\Delta\phi$.
$\cos^{2}\theta_{0}+\sin^{2}\theta_{0}\cos(\phi_{1}-\phi_{2})\leq 1-\frac{2}{\cosh^{2}t_{0}}.$ (108)
After simplification and deformation, it can be observed that this ensures that the effective radius of the sphere, $r_{\text{eff}}=\ell_{\rm dS}\cosh t_{0}\sin\theta_{0}\sin\frac{\Delta\phi}{2}$, is less than or equal to $\ell_{\rm dS}$. Therefore, the entanglement entropy is
$S=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{1}{(\cosh t_{0}\sin\theta_{0}\sin\phi_{a})^{2}}-1}}\right).$ (109)
@@
$S=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{1}{\sin^{2}\frac{\Delta\phi}{2}}-1}}\right)=\frac{c_{\rm dS}}{6}\Delta\phi.$ (110)
By replacing the field theory parameter $c_{\rm dS}$ with the gravitational parameter $\frac{3\ell_{\rm dS}}{2G}$, we find that the above result matches the holographic calculation presented in ref. Kawamoto:2023nki , confirming the validity of half-dS holography. By comparing Eqs. (96) and (109), we find that the entanglement entropy results for spatial intervals, calculated independently in dS/dS holography and half-dS holography, reveal a close analogy between the two frameworks. Therefore, the discussion of SSA aligns with that in dS/dS holography. However, due to the difference in boundary-induced metrics, the BSSA receives corresponding corrections. In half-dS holography, the infinitesimal form of the BSSA is
@@
where $\Delta\tilde{\phi}=\sin\theta_{0}\Delta\phi$. Substituting the entanglement entropy result into the above expression, we obtain
$\text{BSSA}(\Delta\phi)=\frac{c_{dS}}{48}\frac{\sin^{2}(\Delta\phi\sin\theta_{0})}{\sin^{6}(\frac{1}{2}\Delta\phi\sin\theta_{0})(\frac{1}{\sin^{2}\theta_{0}\sin^{2}(\frac{1}{2}\Delta\phi\sin\theta_{0})}-1)^{\frac{3}{2}}}>0.$ (112)
The result is illustrated in figure 17. Compared with figure 13(b), it is evident that figure 17 exhibits only slight differences. The primary behavior of the function remains nearly identical, and a clear deviation in behavior is observed at $\theta_{0}=\frac{\pi}{2}$, indicating that the boundary dual field theory in half-dS holography continues to exhibit non-locality.
Figure 17: The boosted strong subadditivity is violated in half-dS holography.
Similarly, there is a critical position $\sin\theta_{0}\cosh t_{0}=1$ in half-dS holography, as indicated by the red dot in figure 18. As depicted in figure 18, the orange line segment represents $\sin\theta_{0}\cosh t_{0}>1$, where the SSA is violated, while the blue line segment represents $\sin\theta_{0}\cosh t_{0}<1$, where the SSA is satisfied, but the BSSA is still violated.
Figure 18: Half-de Sitter holography. The dual field theory resides on a timelike boundary at $\theta=\theta_{0}$. The blue intervals indicate $\cosh t\leq\frac{1}{\sin\theta_{0}}$, i.e., $\theta_{0}-\frac{\pi}{2}\leq T\leq\frac{\pi}{2}-\theta_{0}$.
We have computed the entanglement entropy of subsystem $A$ using the generalized spherical replica method within the dS/dS holography and half-dS holography frameworks. By verifying the SSA and BSSA, we have found that the entanglement entropy exhibits similar behavior in these two holographic models and that the boundary dual field theory exhibits non-locality. Comparing figures 16 and 18, we observe the same critical point, namely, when the region where the dual field theory is located crosses the cosmological horizon in three-dimensional dS3 spacetime, the entanglement entropy violates the SSA, and even within the cosmological horizon, the entanglement entropy still violates the BSSA.
@@ 6 Conclusion and Discussion
This paper systematically investigated the impact of the $T\bar{T}$ deformation on various holographic dualities, including AdS/CFT, dS/CFT, dS/dS, and half-dS holography, by analyzing the entanglement entropy of the single general spatial subinterval in the boundary dual field theory. When the boundary is spherical, and the subsystem is a semicircular arc, the entanglement entropy can be derived from the spherical partition function on the $n$-replica manifold. For general boundary manifolds, we first mapped it onto a conformal spherical manifold via CHM mapping. Then, we calculated the entanglement entropy using the corresponding effective spherical radius, and derived the correction after $T\bar{T}$ deformation in both the dS Poincaré spacetime and global dS spacetime. In addition, we found that the dual field theory exhibits notable similarities in properties between the dS/dS holography and half-dS holography models. Specifically, the dual field theory in both cases exhibited non-local characteristics. When the dual field theory lies within the cosmological horizon, it satisfies strong subadditivity but violates boosted strong subadditivity. Conversely, when the dual field theory lies outside the cosmological horizon, both strong subadditivity and boosted strong subadditivity are violated.
@@ Acknowledgements.
We would like to thank Shan-Ming Ruan, Yuan Sun for reading our draft and providing comments. We also thank Yang He, and Shan-Ping Wu for their helpful discussion. This work was supported by the National Natural Science Foundation of China (Grants No. 12475056, No. 12235016, No. 12075101, No. 12475053, and No. 12247101), the 111 Project under (Grant No. B20063), the Major Science and Technology Projects of Gansu Province, and Lanzhou City’s scientific research funding subsidy to Lanzhou University. S.H. would appreciate the financial support from the Fundamental Research Funds for the Central Universities and Max Planck Partner Group and the Natural Science Foundation of China (NSFC) Grants No. 12075101 and No. 12235016. L.Z. is supported by the Science and Technology Development Plan Project of Jilin Province, China (No. 20240101326JC).
7 Data Availability Statement
The data that support the findings of this article are openly available ChangHe2025Data
Appendix A Integration over the endpoints of the entanglement region
For a two-dimensional spherical manifold $\mathcal{M}$, on the slice where $\phi=\text{constant}$, we select $\theta\in[0,\pi]$ as the subregion $A$, the entanglement entropy $S_{A}$ can be obtained through the $C$-function. Differentiating $S_{A}$ with respect to $r$, we have
$=2\pi r^{2}\lim_{n\rightarrow 1}n^{2}\int_{0}^{\pi}\sin\theta d\theta\,\partial_{n}\text{tr}\,T.$ (113)
Solving the trace flow and conservation equations without assuming spherical symmetry must determine the energy-momentum tensor $T_{ab}$. For the $n$-replica sphere (27), the solution for the energy-momentum tensor is known Donnelly:2018bef and is given by
A.1 plane
For a plane manifold $\mathcal{M}$, we first map $\mathcal{M}$ to $\mathbb{S}^{2}$ by CHM map, ensuring that the image of the subsystem $A$ under this mapping is a meridian on $\mathbb{S}^{2}$. The mapped metric is conformally equivalent to the sphere, denoted as $\Omega\mathbb{S}^{2}$. After applying the $n$-replica trick, we assume the metric takes the form [^8] We use capital Greek letters to represent the angular coordinates of the conformal sphere $\Omega\mathbb{S}^{2}$
where the conformal factor $\Omega(\Theta,\Phi)$ is constant at the branch points $\Theta=0$ and $\Theta=\pi$, denoted as $\Omega_{0}$ and $\Omega_{\pi}$, respectively. For this metric, eqs. (29) and (30) remain valid. Thus, we have
$r\frac{d}{dr}S(r)$ $=-\lim_{n\rightarrow 1}(1-n\partial_{n})nr^{2}\int_{\Omega\mathbb{S}_{n}^{2}}\Omega^{2}(\Theta,\Phi)\sin\Theta\,d\Theta d\Phi\,\text{tr}\,T$
Since the conformal factor $\Omega(\Theta,\Phi)$ depends on the azimuthal angle $\Phi$, $T_{ab}$ cannot be assumed to depend solely on $\Theta$. Therefore, the $\Phi$ coordinate is not integrated. As discussed in ref. Lewkowycz:2019xse , for the partition function on the sphere (113), in the limit $n\rightarrow 1$, the main contribution to the integral (113) comes from a small neighborhood near the branch points, with radius $\epsilon\sim\sqrt{n-1}$. This property is expected to hold for any manifold $\mathcal{M}$, so the integral (118) is also performed only in small neighborhoods near the branch points. In these regions, the conformal factor $\Omega$ is constant, and the metric (117) can be approximated by the $n$-replica sphere with radius $r\Omega$. As a result, the conformal transformation can be regarded as an identity transformation999 In standard CFT calculations, a conformal transformation is often applied to map $\Omega\mathbb{S}^{2}$ to the spherical metric, and conformal symmetry ensures that the results for $\Omega\mathbb{S}^{2}$ and $\mathbb{S}^{2}$ differ only by a function of the conformal factor. However, for the $T\bar{T}$-deformed field theory, conformal symmetry is absent, and the relationship between quantities before and after the conformal transformation is not determined.. Therefore, partition functions on any manifold $\mathcal{M}$ can be calculated using the results on $\mathbb{S}^{2}$ for non-conformal field theories. Finally, we can get
$rS^{\prime}(r)=\lim_{n\rightarrow 1}2\pi(r\Omega)^{2}\left(\int_{0}^{\Theta_{0}}+\int_{\pi-\Theta_{0}}^{\pi}\right)\sin\Theta\,d\Theta\,\partial_{n}\text{tr}\,T,$ (119)
where $\text{tr}\,T$ corresponds to the solution on the sphere with radius $r\Omega$. When Poincaré coordinates are chosen in EAdS3, the holographic boundary corresponds to a two-dimensional complex plane with the metric
$ds_{2}^{2}=\frac{\ell_{\rm AdS}^{2}}{z_{c}^{2}}(dT^{2}+dX^{2}).$ (120)
The subregion $A$ is defined by $T=0$ and $X\in[-X_{a},X_{a}]$. To compute the entanglement entropy for $A$, the complex plane is mapped onto a two-dimensional sphere such that the image of $A$ corresponds to the meridian on the sphere at $\Phi=0$. This mapping, known as the CHM mapping, is given by
which is conformally equivalent to a sphere with an effective radius $r_{\mathrm{eff}}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$, and the conformal factor $\Omega(\Theta,\Phi)=(1+\cos\Phi\sin\Theta)^{-1}$. As previously noted, when computing the partition function on a general manifold $\mathcal{M}$, the primary contribution to the integral (113) comes from an infinitesimal neighborhood near the branch points. In these regions, the conformal factor $\Omega(\Theta,\Phi)$ can be approximated as 1, implying that the metric (122) effectively describes a sphere with radius
$r_{\mathrm{eff}}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}\,.$ (123)
Thus, within the infinitesimal neighborhoods around the endpoints, transforming (120) to the sphere involves only a conformal transformation with a conformal factor of 1, effectively an identity transformation. Consequently, the partition function remains the same as that of the sphere, with the radius $r$ replaced by the effective radius $r_{\mathrm{eff}}$. Therefore, the solution for the energy-momentum tensor on the $n$-replica sphere is
$T^{\Phi}_{\Phi}$ $=\frac{1}{\pi\lambda_{\rm AdS}}\left(1-\frac{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}}{\sqrt{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\Theta}}}\right)\,.$ (124)
Substituting this into eq. (113) and performing the integration near the endpoints yields the $C$-function as
So we can get the entanglement entropy of the subregion A as
$S_{A}=\frac{c_{\rm AdS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{z_{c}^{2}}{X_{a}^{2}}}}\right)\simeq\frac{c_{\mathrm{AdS}}}{3}\log\frac{2X_{a}}{z_{c}}+\frac{c_{\mathrm{AdS}}}{12}\frac{z_{c}^{2}}{X_{a}^{2}}.$ (126)
Appendix B The gravity calculation
B.1 The dS Poincaré spacetime
@@
Finally we can get that the geodesic distance is
$D_{12}=\ell_{\rm dS}\cos^{-1}\left(\ell_{\rm dS}^{2}\left(1-\frac{2X_{a}^{2}}{\eta_{c}^{2}}\right)\right)\simeq-2i\ell_{\rm dS}\log\frac{2X_{a}}{\eta_{c}}+i\ell_{\rm dS}\frac{\eta_{c}^{2}}{2X_{a}^{2}}+\pi\ell_{\rm dS}.$ (132)
@@ B.2 The global dS spacetime
Appendix C Strong subadditivity and boosted strong subadditivity
For a unitary and Poincaré-invariant field theory, the strong subadditivity is given by:
$S(A)+S(B)-S(A\cup B)-S(A\cap B)\geq 0,$ (139)
where the subsystems $A$ and $B$ are shown in figure 19. This inequality can be derived from the monotonicity of relative entropy, which comes from the locality of field theory Witten:2018zxz . According to the RT formula, the entanglement entropy is quantified by the area of the extremal surface, and thus the entanglement entropy can be expressed as a function of the subsystem size, such as $S(A)\equiv S(L_{A})$. If we set $L_{A\cap B}=L$, $L_{A\cup B}=L+2\delta L$, the lengths of $A$ and $B$ are both $L_{A}=L_{B}=L+\delta L$. In the limit $\delta L\to 0$, the strong subadditivity can be expressed as
$L^{2}S^{\prime\prime}(L)\leq 0,$ (140)
where $L$ represents the size of the subsystems.
Figure 19: 19 Strong subadditivity. 19 Boosted strong subadditivity.
A stronger criterion for proving locality, compared to the strong subadditivity, is the boosted strong subadditivity, which can be derived by applying a boost transformation to subsystems $A$ and $B$, as shown in figure 19. We also set $L_{A\cap B}=L$, $L_{A\cup B}=L+2\delta L$. In Minkowski spacetime, the lengths of $A$ and $B$ are both $L_{A}=L_{B}=\sqrt{L(L+2\delta L)}$. In the limit $\delta L\to 0$, the boosted strong subadditivity can be expressed as
$S^{\prime\prime}(L)+\frac{S^{\prime}(L)}{L}\leq 0.$ (141)